Translate the problem to a system of equations and solve for costs.
Six sandwiches and five sodas cost $36.95. Five sandwiches and five sodas cost $32.20. What is the cost of one sandwich and the cost of one soda?

Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

Six sandwiches and five sodas cost $36.95. Five sandwiches and five sodas cost $32.20. What is the cost of one sandwich and the cost of one soda?
----
Equations:
6d + 5s = 3695 cents
4d + 5s = 3230 cents
-----------------------
Subtract and solve for "d":
2d = 465 cents
d = 232.5 cents
----
A sandwich costs $2.33
----
Substitute and solve for "s":
4(232.5) + 5s = 3230
930 + 5s = 3230
5s = 2300
s = $4.60 (Price of a soda)
================================
Cheers,
Stan H.

Answer by checkley79(3341)

6SA+5SO=36.95
5SA+5SO=32.20 SUBTRACT
---------------------
SA=$4.75 COST OF 1 SANDWICH.
6*4.75+5SO=36.95
28.50+5SO=36.95
5SO=36.95-28.50
5SO=8.45
SO=8.45/5
SO=$1.69 COST OF 1 SODA.
PROOF:
5*4.75+5*1.69=32.20
23.75+8.45=32.20
32.20=32,20